OpenStudy (anonymous):

hat 1 contains three green and four blue tickets. hat 2 contains four green and three blue tickets. one ticket is randomely selected from each hat.

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

i am enjoined from from posting my favorite picture of karnac, but there is no question here. i bet it is going to be "what is the probability of ..."

OpenStudy (anonymous):

oh sorry! what is the probability that the tickets are the same color?

OpenStudy (anonymous):

oops that was wrong probabililty they are both green is \(\frac{3}{7}\times \frac{4}{7}\) and probability they are both blue is \(\frac{4}{7}\times \frac{3}{7}\) multiply and add to get your answer